# Analysis of the strong decay $X(4140)\rightarrow J/\psi\phi$ via the light-cone QCD sum rules

### v1

@@ Abstract
In this article, we take the $X(4140)$ as the axialvector tetraquark state with the symbolic quark structure $sc]_{S}[\bar{s}\bar{c}]_{A}+[sc]_{A}[\bar{s}\bar{c}]_{S}$, and calculate the width of the two-body strong decay $X(4140)\rightarrow J/\psi\phi$ within the framework of the light-cone sum rules. Different from the traditional light-cone sum rules, at the phenomenological side, we introduce parameters $C$ to eliminate the contaminations from the higher resonances and continuum states, and match the hadron side with the QCD side of the correlation function based on rigorous quark-hadron duality to obtain the stable QCD sum rules. Then we obtain the decay width $\Gamma(X(4140)\rightarrow J/\psi\phi)=145\pm 21\,\text{MeV}$, which is in excellent agreement with the experimental data $162\pm 21^{+24}_{-49}\,\text{MeV}$ from the LHCb collaboration. The numerical result supports the possibility that the $X(4140)$ could be the $sc]_{S}[\bar{s}\bar{c}]_{A}+[sc]_{A}[\bar{s}\bar{c}]_{S}$ type axialvector tetraquark state.
@@ 2 Light-cone sum rules for the $X(4140)$
where $P^{\alpha\beta}=V_{\tau^{\prime}}q_{\lambda^{\prime}}\varepsilon^{\lambda^{\prime}\tau^{\prime}\beta\alpha}$, $V^{2}=1$, $V\cdot q=0$ and $V\cdot p=0$. Thereafter, we would like to add a subscript $H$ to represent the hadron side, i.e. $\Pi_{H}(p^{\prime 2},q^{2})$.
@@
Then we set $p^{\prime 2}=q^{2}$ and perform the Borel transformation with respect to the variable $Q^{2}=-q^{2}$ to obtain the light-cone QCD sum rules,
@@ 3 Numerical results and discussions
where $p(a,b,c)=\frac{\sqrt{[a^{2}-(b+c)^{2}][a^{2}-(b-c)^{2}]}}{2a}$. The width $\Gamma_{X(4140)\rightarrow J/\psi\phi}=145\pm 21\,\text{MeV}$ is in very good agreement with the measured data $\Gamma_{X(4140)\rightarrow J/\psi\phi}=162\pm 21^{+24}_{-49}\text{MeV}$ by the LHCb collaboration in 2021. We obtain additional support that the predictions from the QCD sum rules favors assigning the $X(4140)$ as the $sc]_{S}[\bar{s}\bar{c}]_{A}+[sc]_{A}[\bar{s}\bar{c}]_{S}$ type axialvector tetraquark state with $J^{PC}=1^{++}$. We can extend present work to study the strong decays of other tetraquark candidates with the light-cone QCD sum rules to diagnose their nature.
@@ 4 Conclusion
In this article, we take the $X(4140)$ as the 1S $sc]_{S}[\bar{s}\bar{c}]_{A}+[sc]_{A}[\bar{s}\bar{c}]_{S}$ tetraquark state with the $J^{PC}=1^{++}$ to study the hadronic coupling constant $g_{XJ/\psi\phi}$ via the light-cone QCD sum rules, then calculate the width of the two-body strong decay $X(4140)\rightarrow J/\psi\phi$. In calculations, we introduce free parameters to parameterize the contributions of the higher resonances and continuum states, resort to the rigorous quark-hadron duality to match the hadron side with the QCD side of the correlation function, then obtain the light-cone QCD sum rules for the hadronic coupling constant. Through varying the free parameter, we obtain the flat Borel platform and extract the value of the hadronic coupling constant. The corresponding decay width $\Gamma(X(4140)\rightarrow J/\psi\phi)=145\pm 21\,\text{MeV}$ is in very good agreement with the updated measurement $162\pm 21^{+24}_{-49}\,\text{MeV}$ by the LHCb collaboration. The numerical results lead to additional support for the possibility that $X(4140)$ could be the $sc]_{S}[\bar{s}\bar{c}]_{A}+[sc]_{A}[\bar{s}\bar{c}]_{S}$ type axialvector tetraquark state, as we cannot assign a hadron unambiguously with the mass alone.

### v2

@@ Abstract
In this article, we take the $X(4140)$ as the axialvector tetraquark state with the symbolic quark structure $sc]_{S}[\bar{s}\bar{c}]_{A}+[sc]_{A}[\bar{s}\bar{c}]_{S}$, and calculate the width of the two-body strong decay $X(4140)\rightarrow J/\psi\phi$ within the framework of the light-cone sum rules. Different from the traditional light-cone sum rules, at the phenomenological side, we introduce parameters $C$ to eliminate the contaminations from the higher resonances and continuum states, and match the hadron side with the QCD side of the correlation function based on rigorous quark-hadron duality to obtain the stable QCD sum rules. Then we obtain the decay width $\Gamma(X(4140)\rightarrow J/\psi\phi)=145\pm 21\,\text{MeV}$, which is reasonable according to the experimental data $162\pm 21^{+24}_{-49}\,\text{MeV}$ from the LHCb collaboration. The numerical result supports the possibility that the $X(4140)$ could be the $sc]_{S}[\bar{s}\bar{c}]_{A}+[sc]_{A}[\bar{s}\bar{c}]_{S}$ type axialvector tetraquark state.
@@ 2 Light-cone sum rules for the $X(4140)$
where $P^{\alpha\beta}=V_{\tau^{\prime}}q_{\lambda^{\prime}}\varepsilon^{\lambda^{\prime}\tau^{\prime}\beta\alpha}$, the unit vector $V_{\tau}$ is defined by us, which satisfies the conditions: $V^{2}=1$, $V\cdot q=0$ and $V\cdot p=0$, but it does not have a specific expression. Thereafter, we would like to add a subscript $H$ to represent the hadron side, i.e. $\Pi_{H}(p^{\prime 2},q^{2})$.
@@
The spectral densities $\rho_{H}(s^{\prime},s)$ and $\rho_{QCD}(p^{\prime 2},s)$ are physical, while the variables $p^{\prime 2}$ and $q^{2}$ in Eq. (24) are free variables after performing the operator product expansion. Generally speaking, we can set $p^{\prime 2}=\alpha q^{2}$ with $\alpha$ to be a finite quantity. According to the mass poles at $s^{\prime}=m^{2}_{X}$ and $s=m_{J/\psi}^{2}$, we can obtain an approximated relation $s^{\prime}=s$, therefore, we set $p^{\prime 2}=q^{2}$ and perform the Borel transformation with respect to the variable $Q^{2}=-q^{2}$ to obtain the light-cone QCD sum rules,
@@ 3 Numerical results and discussions
where $p(a,b,c)=\frac{\sqrt{[a^{2}-(b+c)^{2}][a^{2}-(b-c)^{2}]}}{2a}$. The width $\Gamma_{X(4140)\rightarrow J/\psi\phi}=145\pm 21\,\text{MeV}$ is reasonable according to the measured data $162\pm 21^{+24}_{-49}\text{MeV}$ by the LHCb collaboration in 2021. We obtain additional support that the predictions from the QCD sum rules favors assigning the $X(4140)$ as the $sc]_{S}[\bar{s}\bar{c}]_{A}+[sc]_{A}[\bar{s}\bar{c}]_{S}$ type axialvector tetraquark state with $J^{PC}=1^{++}$. We can extend present work to study the strong decays of other tetraquark candidates with the light-cone QCD sum rules to diagnose their nature.
@@ 4 Conclusion
In this article, we take the $X(4140)$ as the 1S $sc]_{S}[\bar{s}\bar{c}]_{A}+[sc]_{A}[\bar{s}\bar{c}]_{S}$ tetraquark state with the $J^{PC}=1^{++}$ to study the hadronic coupling constant $g_{XJ/\psi\phi}$ via the light-cone QCD sum rules, then calculate the width of the two-body strong decay $X(4140)\rightarrow J/\psi\phi$. In calculations, we introduce free parameters to parameterize the contributions of the higher resonances and continuum states, resort to the rigorous quark-hadron duality to match the hadron side with the QCD side of the correlation function, then obtain the light-cone QCD sum rules for the hadronic coupling constant. Through varying the free parameter, we obtain the flat Borel platform and extract the value of the hadronic coupling constant. The corresponding decay width $\Gamma(X(4140)\rightarrow J/\psi\phi)=145\pm 21\,\text{MeV}$ is reasonable according to the updated measurement $162\pm 21^{+24}_{-49}\,\text{MeV}$ by the LHCb collaboration. The numerical results lead to additional support for the possibility that $X(4140)$ could be the $sc]_{S}[\bar{s}\bar{c}]_{A}+[sc]_{A}[\bar{s}\bar{c}]_{S}$ type axialvector tetraquark state, as we cannot assign a hadron unambiguously with the mass alone.